OpenStudy (r_josh10):

What is the percent of increase from 60 to 75? A. 80% B. 25% C. 20% D. 15% Please explain and help. Will Fan and Medal.

OpenStudy (r_josh10):

Is the answer 25?

OpenStudy (theedwardsfamily):

i think so. you may be right

563blackghost (563blackghost):

you are correct it is 25% because Increase = 75 - 60 = 15% increase = 15/60 x 100 = 25% so there is a 25% increase

OpenStudy (legends):

No guys what is 75-60

OpenStudy (legends):

I thought it was 15 percent

OpenStudy (r_josh10):

Thankyou @563blackghost I will medal you

563blackghost (563blackghost):

@Legends you just subtracted to get a percent you need to divide by 100

563blackghost (563blackghost):

i mean multiply sorry

563blackghost (563blackghost):

And your welcome @r_josh10